OpenStudy (anonymous):

how do I find a quadratic model for a set of values? (-1,1) , (1,1) (3,9)

jimthompson5910 (jim_thompson5910):

The general quadratic equation is y = a*x^2 + b*x + c We'll use this to set up our equations

jimthompson5910 (jim_thompson5910):

The first point is (-1,1). So x = -1 and y = 1 y = a*x^2 + b*x + c y = a*(-1)^2 + b*(-1) + c ... plug in x = -1 y = a*(1) + b*(-1) + c y = 1a - 1b + c 1 = 1a - 1b + c ... plug in y = 1 1 = a - b + c a - b + c = 1 So the first equation is `a - b + c = 1`

jimthompson5910 (jim_thompson5910):

The second point is (1,1). So x = 1 and y = 1 y = a*x^2 + b*x + c y = a*(1)^2 + b*(1) + c ... plug in x = 1 y = a*(1) + b*(1) + c y = 1a + 1b + c 1 = a + b + c ... plug in y = 1 a + b + c = 1 The second equation is `a + b + c = 1`

jimthompson5910 (jim_thompson5910):

And finally, the third point is (3,9), so... x = 3 y = 9 y = a*x^2 + b*x + c y = a*(3)^2 + b*(3) + c ... replace x with 3 y = 9a + 3b + c 9 = 9a + 3b + c ... replace y with 9 9a + 3b + c = 9 The third equation is `9a + 3b + c = 9`

jimthompson5910 (jim_thompson5910):

we now have this system of equations \[\Large \begin{cases}a-b+c = 1\\a+b+c = 1\\9a+3b+c = 9\end{cases}\] hopefully you're able to solve for a,b,c ?

OpenStudy (anonymous):

yep! thanks alot :D

jimthompson5910 (jim_thompson5910):

you're welcome